The measure of each of the base angles of an isosceles triangle is 24 degrees less than the measure of the vertex angle. Find the measure of the vertex angle.
Answer by mathitutor(25)

let the vertex angle = x
Hence the each of base angle = x - 24
Sum of all angles of a triangle = 180
=> x + x - 24 + x - 24 = 180
=> 3x - 48 = 180
=> 3x = 180 + 48
=> 3x = 228
=> x = 76 degree.